OpenStudy (countrygirl1431):

The volume of a box(V) varies directly with its length(l). If one of the boxes has a volume of 288 cubic inches and a length of 12 inches, what is the constant of proportionality for the group of boxes? 3,456 24 1/24

OpenStudy (countrygirl1431):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm going to represent the length as upper case L instead of lower case l

jimthompson5910 (jim_thompson5910):

The volume of a box(V) varies directly with its length(L). so that means V = k*L where k is some fixed number

OpenStudy (countrygirl1431):

ok

jimthompson5910 (jim_thompson5910):

`one of the boxes has a volume of 288 cubic inches and a length of 12 inches` so... V = 288 L = 12 plug these values into the equation V = k*L to get... V = k*L 288 = k*12 now solve for k to get k = ???

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (countrygirl1431):

yes

OpenStudy (countrygirl1431):

@jim_thompson5910 would i divide both sides by 12 to get my answer

jimthompson5910 (jim_thompson5910):

yes you would

jimthompson5910 (jim_thompson5910):

to undo the multiplication of 12

OpenStudy (countrygirl1431):

k=24

jimthompson5910 (jim_thompson5910):

correct